# Temporal Conformal Prediction (TCP): A Distribution-Free Statistical and Machine Learning Framework for Adaptive Risk Forecasting

### v4

@@ 6.  RESULTS AND DISCUSSION
Implementation note (comparability). All learning-based methods use the same feature set. TCP/TCP-RM are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points. Rolling QR is also evaluated out-of-sample using a 192-point trailing training window. The slight differences in the “Predictions” counts (e.g., 1508 for QR vs. 1448 for TCP/TCP-RM) reflect different warm-up requirements (QR does not reserve a calibration slice). Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: QR is sharp but under-calibrated; TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
6.1 Hyperparameter Sensitivity (SP500, TCP-RM)
We stress–test the adaptive layer by varying the rolling window $w$ and the initial step size $\gamma_{0}$ for TCP-RM(Table 2). For comparability we fix the calibration slice at $|\mathcal{C}|=40$ and sweep $w\in\{100,252,500\}$ and $\gamma_{0}\in\{0.005,0.01,0.05\}$ on SP500. Results for BTC-USD and Gold exhibit the same pattern and are reported in Appendix D.
Hyperparameters Performance
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9575 5.3052
100 0.010 0.9575 5.3066
100 0.050 0.9581 5.3172
252 0.005 0.9579 5.1168
252 0.010 0.9579 5.1191
252 0.050 0.9586 5.1314
500 0.005 0.9525 5.4135
500 0.010 0.9525 5.4319
500 0.050 0.9567 5.5348
What we learn. (i) Across these settings, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), while w=100w{=}100 and w=500w{=}500 are wider ($\approx 5.31$ and 5.415.415.535.53, respectively), with coverage remaining near or just above the 95% target for $w\in\{100,252\}$ and slightly lower for some w=500w{=}500 configurations. (ii) At a fixed $w$, increasing $\gamma_{0}$ produces only minor changes: coverage nudges up by at most a few tenths of a percentage point with a negligible width increase (e.g., at w=252w{=}252, $\gamma_{0}{=}0.050$ lifts coverage from $0.9579$ to $0.9586$ with a $\sim 0.012$ increase in width; at w=500w{=}500, $\gamma_{0}{=}0.050$ raises coverage from $0.9525$ to $0.9567$ while widening intervals modestly).
Takeaway. TCP–RM is robust over $(w,\gamma_{0})$ in this range. For the main experiments, w=252w{=}252 with $\gamma_{0}{=}0.01$ offers a good balance (coverage $\approx 0.9579$, width $\approx 5.1191$). If one prefers higher coverage at w=500w{=}500, a slightly larger $\gamma_{0}$ (e.g., $0.05$) helps, at the cost of modestly wider bands.
We also perform sensitivity analysis on TCP. The results are given in Appendix E.
6.2 Visualization for TCP vs. TCP-RM on S&P 500 (95% intervals)
In this section, we present the TCP vs TCP-RM trace plot for S&P 500.
Figure 2: TCP vs. TCP-RM on S&P 500 (95% intervals). Top: daily returns with two‐sided prediction bands for TCP (blue) and TCP-RM (purple). Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target. Bottom: evolution of conformal thresholds: split‐conformal level $C_{\text{split}}$, online Robbins–Monro offset $C_{\text{RM}}$, and the effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$. During the COVID-19 crash the bands widen sharply, the rolling coverage briefly departs from the 95% target and then re-enters the band, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
In Fig. 2, TCP and TCP-RM behave similarly in tranquil periods, but during stress the effective threshold $C_{\text{eff}}$ (bottom panel) rises quickly, producing wider bands (top) and steering the 30-day rolling coverage back toward the 95% target (middle). The Robbins–Monro offset $C_{\text{RM}}$ remains near zero except around regime shifts, where it provides the extra adjustment that closes the gap to target coverage. This visual evidence complements Table 1: on S&P 500, TCP-RM attains $\approx$95% coverage with only a marginal width increase relative to TCP.
@@ 7.  CONCLUSION AND FUTURE WORK
This paper introduced Temporal Conformal Prediction (TCP), a simple and effective framework for calibrated prediction intervals in nonstationary time series. TCP combines a rolling, exchangeability‐aware split–conformal layer with a modern quantile forecaster, and the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
Across three asset classes, TCP/TCP–RM delivers a strong calibration–sharpness balance: coverage is close to the 95% target; in our evaluation, the intervals are wider than Historical Simulation, reflecting stricter calibration during volatile periods. In contrast, the rolling QR baseline attains the narrowest bands but undershoots the target coverage, especially in turbulent regimes. Crisis‐window visualizations around the COVID‐19 crash illustrate TCP’s intended adaptiveness: intervals widen rapidly as volatility spikes and contract as conditions stabilize, with red dots indicating miscoverage events. A targeted sensitivity analysis shows robustness to reasonable choices of rolling window and Robbins–Monro step size, and suggests default settings that are stable in practice.
Limitations. Like any sequential method, TCP can temporarily under‐cover during abrupt regime shifts before its calibration catches up. Performance also inherits the quality of the base quantile forecaster and the choice of windowing; extremely heavy tails or structural breaks can stress any fixed feature set. Finally, while TCP–RM improves calibration at negligible width cost on average, very aggressive step sizes can overcorrect.
Practical implications. By providing distribution‐free intervals that adapt in time, TCP aligns naturally with backtesting and governance needs in risk management (coverage targets, exception tracking, and scenario robustness). The framework is lightweight (single threshold) and model‐agnostic (plugs into any quantile forecaster), making deployment straightforward in production pipelines.
7.1 Future Work
Several extensions follow naturally:
  1. 1.
    Multivariate and portfolio TCP. Generalize to joint/conditional quantiles for baskets of assets; couple TCP with copula or factor models to control portfolio‐level coverage.
  2. 2.
    Conditional and localized coverage. Target coverage conditional on covariates (e.g., volatility state) and explore adaptive nonconformity scores that emphasize tail risk or signed errors.
  3. 3.
    Richer forecasters and features. Swap the base learner for neural quantile models or boosted variants with realized‐volatility inputs (intraday RV), order‐flow, or macro features; assess benefits for sharpness at fixed coverage.
  4. 4.
    Change–point aware adaptation. Trigger temporary step–size boosts or threshold projections when a regime break is detected to shorten the transient under/over‐shoot in coverage.
  5. 5.
    Heavy‐tail robustness. Investigate EVT‐inspired nonconformity scores or asymmetric penalties to better stabilize performance under extreme shocks.
  6. 6.
    Decision integration. Map TCP intervals to VaR/ES, margin, and capital rules; evaluate utility/capital efficiency vs. conservative baselines in realistic backtests.
  7. 7.
    Beyond finance. TCP’s adaptiveness is broadly applicable wherever nonstationarity and calibrated uncertainty matter:
    • Biostatistics/medicine: ICU vitals monitoring and glucose‐range forecasting with patient–specific adaptation; adaptive dosing intervals under drift.
    • Epidemiology: Calibrated intervals for nowcasts/forecasts of incidence and $R_{t}$ that widen under reporting delays or testing-policy shifts, plus adaptive risk bands for hospitalization/ICU load as transmission regimes change.
    • Genetics/Genomics: Uncertainty bands for polygenic risk scores and eQTL effect sizes that remain calibrated under population/covariate shift; interval forecasts for gene expression (bulk or single-cell) with batch effects and sequencing-depth drift; calibrated pathogenicity scores for rare variants.
    • Sports analytics: Game-state–aware intervals for win probability, expected goals/points, and player performance that adapt to tempo and lineup changes; injury-risk ranges for load management with rapid recalibration after shocks (e.g., in-game injuries or schedule congestion).
    • Renewable energy: Wind/solar generation intervals that expand/contract with weather regimes to aid grid balancing.
    • Operations/marketing: Demand and lead‐time intervals for inventory control; calibrated lift ranges in marketing‐mix models under campaign drift.
    • Transportation and climate: Travel‐time intervals under traffic regime shifts; local hazard/risk bands for extreme weather.
@@ 7.1 Future Work
In sum, TCP offers a compact, distribution‐free, and adaptive route to calibrated uncertainty quantification in the presence of temporal drift. Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
@@ Appendix A MATHEMATICAL PROOFS
Appendix B SUPPLEMENTARY VISUALIZATIONS
@@
Figure 3: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 4: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist) for Gold daily returns during the COVID-19 market crash (February-April 2020).
Appendix C VISUALIZATIONS OF THE TCP VS TCP-RM TRACE PLOTS
This section provides the visualizations of the TCP vs TCP-RM trace plots for BTC-USD (Figure 5) and Gold (Figure 6), complementing the main visualization for the S&P 500 presented in the results section.
Figure 5: TCP vs. TCP-RM on BTC-USD (95% intervals). Same layout as Fig. 2. Large, frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, keeping rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
Figure 6: TCP vs. TCP-RM on Gold (95% intervals). Same layout as Fig. 2. Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; elsewhere, $C_{\text{RM}}$ stays near zero and the method behaves like standard split-conformal TCP.
Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
This section provides the full sensitivity analysis results for BTC-USD (Table 3) and Gold (Table 4), complementing the main analysis on the S&P 500 presented in the results section. The findings confirm that the TCP-RM framework is robust and performs consistently across different asset classes and hyperparameter settings. For BTC-USD, widths decrease as $w$ grows (tightest at w=500w{=}500) while coverage remains near 95%–96%; $\gamma_{0}$ has negligible impact within a fixed $w$. For Gold, widths are smallest at w=100w{=}100 and increase with $w$; coverage hovers around the nominal level; $\gamma_{0}$ effects are minor at fixed $w$.
@@ Appendix D SUPPLEMENTARY SENSITIVITY ANALYSIS
Appendix E SUPPLEMENTARY SENSITIVITY ANALYSIS ON TCP (NO RM)
For completeness, in Tables 5, 6 and 7, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.

### v5

@@ 6.  RESULTS AND DISCUSSION
Implementation note (comparability). All learning-based methods (TCP, TCP-RM, ACI, QR, QR-Linear) use the same feature set. TCP/TCP-RM and ACI are evaluated out-of-sample with a rolling window of 252 points split into 192 training and 60 calibration points. Rolling QR and QR-Linear are also evaluated out-of-sample using a 192-point trailing training window without a separate calibration slice. The differences in the “Predictions” counts (e.g., 1508 for QR/QR-Linear vs. 1448 for TCP/TCP-RM/ACI, 1468 for Hist, 1670 for GARCH) reflect these distinct warm-up requirements. Aligning the evaluation horizon to the longest warm-up yields qualitatively identical conclusions: tree-based QR remains sharp but under-calibrated; QR-Linear and ACI partially trade width for better coverage yet stay sub-nominal; and TCP/TCP-RM are near-nominal and, in this evaluation, wider than Historical Simulation.
6.1 Backtests of tail coverage
To complement marginal coverage and interval width, we perform standard VaR-style backtests on the 5% lower tail using the Kupiec (Kupiec1995) unconditional coverage test (LR_uc), the Christoffersen (Christoffersen1998) independence test (LR_ind), and their joint conditional coverage test (LR_cc). Table 3 reports the likelihood-ratio statistics, $p$-values, the total number of exceedances (days where the realized return falls below the nominal 5% lower bound), and the number of out-of-sample predictions $n$ for the S&P 500.
Table 3: VaR-style backtests on the 5% lower tail for S&P 500.
Model LR_uc p_uc LR_ind p_ind LR_cc p_cc Exceed. $n$
TCP 0.170625 0.679557 2.022260 0.155008 2.192885 0.334057 69 1448
TCP-RM 0.287050 0.592117 2.203339 0.137712 2.490389 0.287885 68 1448
ACI 127.325197 0 7.171877 0.007406 134.497073 0 183 1448
QR 280.342540 0 1.404329 0.236000 281.746869 0 253 1508
QR-Linear 100.761827 0 10.714099 0.001063 111.475925 0 174 1508
GARCH 334.481845 0 0.022234 0.881466 334.504079 0 289 1670
Hist 9.827570 0.001719 22.904448 0.000002 32.732018 0 101 1468
On S&P 500, TCP and TCP-RM are the only methods that pass all three backtests at conventional significance levels: their LR_uc, LR_ind, and LR_cc $p$-values stay well above 0.05, and the number of exceedances (68–69) is close to the 5% target ($\approx 0.05\times 1448$). By contrast, ACI, QR, QR-Linear, GARCH, and Historical Simulation are all strongly rejected by at least one of the tests: their LR_uc and LR_cc statistics are very large with $p$-values numerically equal to zero (e.g., QR: 253 exceedances; GARCH: 289; Hist: 101). This confirms that, on equities, only TCP/TCP-RM deliver statistically acceptable 95% tail coverage. The corresponding backtests for BTC-USD and Gold show the same qualitative pattern and are reported in Appendix D.
We also perform sensitivity analyses for TCP-RM and TCP, with results reported in Appendix F and Appendix G. TCP vs. TCP-RM trace plots are presented in Appendix H.
@@ 7.  CONCLUSION AND FUTURE WORK
This paper introduced Temporal Conformal Prediction (TCP), a simple framework for calibrated prediction intervals in nonstationary time series. TCP combines a rolling, exchangeability-aware split-conformal layer with a modern quantile forecaster, while the TCP–RM variant adds a single Robbins–Monro offset that adjusts the conformal threshold online using observed coverage errors.
Across three asset classes, TCP and TCP–RM achieve a strong calibration–sharpness balance: coverage stays close to the 95% target and the intervals are wider than Historical Simulation, reflecting stricter calibration in volatile periods. The rolling tree-based QR model and the classical linear baseline (QR-Linear) produce the narrowest bands but undershoot the target coverage, especially in turbulent regimes, and the adaptive conformal baseline (ACI) only partially corrects this undercoverage. In our study, TCP and TCP-RM are the only methods that simultaneously passes standard VaR backtests and maintains competitive sharpness on S&P 500, BTC-USD, and Gold.
Like any sequential method, TCP can temporarily under-cover during abrupt regime shifts before calibration catches up. Performance depends on the quality of the base quantile forecaster and the windowing scheme, and very heavy tails or structural breaks can stress any fixed feature set. In our experiments, TCP–RM leaves average calibration and sharpness essentially unchanged at the chosen hyperparameters, but its Robbins–Monro layer provides a principled online mechanism that can correct coverage drift when step sizes are tuned more aggressively; overly aggressive step sizes can, however, overcorrect. On the positive side, TCP provides distribution-free, time-adaptive intervals that align naturally with backtesting and governance needs in risk management, and its single-threshold, model-agnostic design makes deployment straightforward in production pipelines.
Several extensions follow naturally. On the modeling side, it is natural to develop multivariate and portfolio versions of TCP, target conditional or localized coverage (for example by volatility state), explore richer quantile forecasters and feature sets, design change-point-aware adaptations that shorten transients after regime breaks, and study heavy-tail robustness through EVT-inspired or asymmetric nonconformity scores. On the application side, TCP intervals can be mapped directly to VaR/ES, margin, and capital rules and evaluated in realistic backtests for utility and capital efficiency. Beyond finance, the same adaptive, distribution-free mechanism could support calibrated uncertainty in domains such as biostatistics and medicine (for example ICU vitals or glucose-range forecasting), epidemiology, genetics and genomics, sports analytics, renewable energy, and operations or transportation planning wherever nonstationarity and reliable uncertainty quantification are central.
@@
In sum, TCP offers a compact, distribution-free, and adaptive route to calibrated uncertainty quantification under temporal drift. Its empirical performance, visual adaptiveness in crisis windows, and robustness to sensible hyperparameters suggest that it is a practical default for risk forecasting and a promising building block for broader sequential decision systems.
Code Availability
@@ Appendix A MATHEMATICAL PROOFS
Appendix B ADDITIONAL CRISIS-WINDOW RESULTS
Table 4 and Table 5 report the COVID crisis-window coverage and widths for BTC-USD and Gold.
Table 4: Crisis-window empirical coverage and average width of 95% prediction intervals for BTC-USD during COVID (Feb–Apr 2020, $n=55$ days).
Model Coverage Avg. width $n$
TCP 0.8909 24.2814 55
TCP-RM 0.8909 24.2729 55
QR (rolling) 0.8909 17.4407 55
GARCH 0.8545 17.9290 55
Hist 0.9091 19.8499 55
ACI 0.8727 20.2050 55
Table 5: Crisis-window empirical coverage and average width of 95% prediction intervals for Gold during COVID (Feb–Apr 2020, $n=55$ days).
Model Coverage Avg. width $n$
TCP 0.8727 6.8087 55
TCP-RM 0.8727 6.8119 55
QR (rolling) 0.7455 4.0744 55
GARCH 0.8000 4.9112 55
Hist 0.6909 3.9957 55
ACI 0.5818 2.6643 55
Appendix C ADDITIONAL TCP VS TCP-RM CRISIS-WINDOW PLOTS
Figure 3 and Figure 4 show the TCP and TCP-RM 95% prediction intervals alongside returns for BTC-USD and Gold during the COVID crisis window (Feb–Apr 2020).
Figure 3: BTC-USD: TCP vs TCP-RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).
Figure 4: Gold: TCP vs TCP-RM 95% prediction intervals during the COVID crisis window (Feb–Apr 2020).
Appendix D ADDITIONAL BACKTESTS
Table 6 reports the same Kupiec and Christoffersen 5% lower-tail backtests for BTC-USD and Gold. The qualitative pattern matches the S&P 500 results in the main text: TCP and TCP-RM are the only models that are not rejected by any of the three tests on either asset, while ACI, QR, QR-Linear, GARCH, and Historical Simulation accumulate far too many exceptions and fail at least one of the diagnostics.
Table 6: VaR-style backtests on the 5% lower tail for BTC-USD and Gold.
Asset Model LR_uc p_uc LR_ind p_ind LR_cc p_cc Exceed. $n$
BTC-USD TCP 0.434337 0.509868 1.094155 0.295552 1.528492 0.465685 67 1448
TCP-RM 0.434337 0.509868 1.094155 0.295552 1.528492 0.465685 67 1448
ACI 89.465919 0 11.093370 0.000866 100.559289 0 163 1448
QR 192.386083 0 0.868073 0.351489 193.254155 0 218 1508
QR-Linear 73.429698 0 5.932611 0.014863 79.362309 0 158 1508
GARCH 221.469896 0 1.342326 0.246624 222.812222 0 245 1670
Hist 1.023574 0.311673 9.776309 0.001768 10.799882 0.004517 82 1468
Gold TCP 1.563187 0.211199 0.338333 0.560793 1.901519 0.386447 83 1448
TCP-RM 1.563187 0.211199 0.338333 0.560793 1.901519 0.386447 83 1448
ACI 280.191360 0 2.694226 0.100712 282.885586 0 247 1448
QR 288.442776 0 1.452719 0.228092 289.895494 0 256 1508
QR-Linear 78.288019 0 13.470197 0.000242 91.758217 0 161 1508
GARCH 291.399454 0 1.952267 0.162343 293.351721 0 273 1670
Hist 8.513620 0.003525 2.757010 0.096829 11.270630 0.003570 99 1468
Appendix E ADDITIONAL VISUALIZATIONS OF THE PREDICTION INTERVALS
@@
Figure 5: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for BTC-USD daily returns during the COVID-19 market crash (February-April 2020).
Figure 6: A comparison of 95% prediction intervals from five models (TCP, TCP-RM, QR, GARCH, Hist, ACI) for Gold daily returns during the COVID-19 market crash (February-April 2020).
Appendix F SENSITIVITY ANALYSIS FOR TCP-RM
This section reports the full hyperparameter sensitivity analysis for TCP-RM across the S&P 500, BTC-USD, and Gold. For each asset we vary the rolling window size $w\in\{100,252,500\}$ and the initial Robbins-Monro step size $\gamma_{0}\in\{0.005,0.01,0.05\}$ while keeping the calibration slice fixed at $|\mathcal{C}|=40$. For every configuration, we report empirical coverage of the nominal 95% prediction intervals and average interval width (Tables 7,8, and 9).
Hyperparameters Performance
Window Size ($w$) $\gamma_{0}$ Coverage Interval Width
100 0.005 0.9575 5.3052
100 0.010 0.9575 5.3066
100 0.050 0.9581 5.3172
252 0.005 0.9579 5.1168
252 0.010 0.9579 5.1191
252 0.050 0.9586 5.1314
500 0.005 0.9525 5.4135
500 0.010 0.9525 5.4319
500 0.050 0.9567 5.5348
@@ Appendix F SENSITIVITY ANALYSIS FOR TCP-RM
Summary. On S&P 500, w=252w{=}252 yields the narrowest intervals (width $\approx 5.12$), with w=100w{=}100 and w=500w{=}500 producing slightly wider bands and coverage close to or just above the 95% target, except for mild undercoverage in some w=500w{=}500 settings; this motivates our choice w=252w{=}252, $\gamma_{0}{=}0.01$ in the main experiments. For BTC-USD, interval widths shrink as $w$ increases, with the tightest bands at w=500w{=}500, while coverage remains stably around 95–96% across all configurations. For Gold, widths are smallest at w=100w{=}100 and grow with $w$, and coverage stays near the nominal level throughout. Across all three assets, TCP-RM is therefore stable over the $(w,\gamma_{0})$ grid considered here: the window length $w$ controls the sharpness–stability tradeoff, whereas the specific choice of $\gamma_{0}$ within this range has only a minor effect on coverage and width.
Appendix G SENSITIVITY ANALYSIS ON TCP (NO RM)
For completeness, in Tables 10, 11 and 12, we also report sensitivity for TCP (split-conformal only, no Robbins–Monro offset). We vary $w\in\{100,252,500\}$ with $|\mathcal{C}|=40$ for each asset.
@@ Appendix G SENSITIVITY ANALYSIS ON TCP (NO RM)
Appendix H TCP VS. TCP-RM TRACE PLOTS
This section provides TCP and TCP-RM trace plots for all three assets (Figures 7, 8 and 9. Each panel shows daily returns, the corresponding 95% prediction intervals, a rolling coverage diagnostic, and the evolution of the conformal thresholds.
Figure 7: TCP vs. TCP-RM on S&P 500 (95% intervals). Top: daily returns with two-sided prediction bands for TCP (blue) and TCP-RM (purple). Middle: 30-day rolling empirical coverage; the dashed line marks the 95% target. Bottom: evolution of conformal thresholds: split-conformal level $C_{\text{split}}$, Robbins–Monro offset $C_{\text{RM}}$, and effective total $C_{\text{eff}}=C_{\text{split}}+C_{\text{RM}}$. During the COVID-19 crash the bands widen sharply, the rolling coverage briefly moves away from the 95% target and then returns, and $C_{\text{eff}}$ spikes before decaying as conditions stabilize.
Figure 8: TCP vs. TCP-RM on BTC-USD (95% intervals). Same layout as Figure 7. Large and frequent shocks lead to sustained but controlled elevations in $C_{\text{eff}}$, which keeps rolling coverage close to 95% while allowing intervals to shrink as volatility recedes.
Figure 9: TCP vs. TCP-RM on Gold (95% intervals). Same layout as Figure 7. Spikes in $C_{\text{eff}}$ are concentrated around crisis windows; away from those periods $C_{\text{RM}}$ stays close to zero and the method behaves like standard split-conformal TCP.